OpenStudy (anonymous):

Not sure what to use to solve this or where to start. Given the unbalanced equation Na+ O2 > Na20 how many grams of Na2O could be produced from 0.400 moles of Na and excess O2?

OpenStudy (australopithecus):

Na+ O2 -> Na20 split atoms up on both side keeping compounds together ( 1 Na ),( 2 O ) ->( 2 Na, 1 O) try doubling the sodium (2 Na ),( 2 O ) ->( 2 Na, 1 O) we still have an imbalance of oxygen try doubling the Na20 (2 Na ),( 2 O ) ->( 4 Na, 2 O) Now we have an imbalance of sodium again double it again (4 Na ),( 2 O ) ->( 4 Na, 2 O) Ok looks balanced so we have 4Na + O2 -> 4Na2O ok now we have 0.400mol of sodium a ratio of 4 sodium per 4 Na2O produced so lets convert \[0.400mol_{Na}*\frac{4mol_{Na_2O}}{4mol_{Na}} = 0.400mol_{Na_2O}\] so we see that they are 1 to 1 (but this is a good way to convert) so we get 0.400mol of Na2O now to convert to grams we use the formula \[Moles = \frac{Grams}{MolecularMass}\] the molecular mass of Na2O is equal to 22.99g/mol + 22.99g/mol + 15.9994g/mol = 61.9789g/mol so 0.400mol = x/61.9789g/mol 0.400mol*61.9789g/mol = x x = 24.79g

OpenStudy (australopithecus):

hope that helps

OpenStudy (anonymous):

it helps a bit, but there are still things i'm confused about. I don't quite understand how this (0.400molNa∗4molNa2O4molNa=0.400molNa2O) conversion worked. forgive me i'm horrible at math. Also, the answer you got didn't match up with the available answers for me to circle, which confuses me because your answer makes sense! the available answers are a.) 6.20, b.) 9.20, c.) 3.10, d.) 4.60, and e.) 12.4.

OpenStudy (australopithecus):

Alright I will check it out right away I'm prone to mistakes but do you mean you do not understand gravimeter factor?

OpenStudy (australopithecus):

Sorry I mean gravimetric factor

OpenStudy (australopithecus):

Well I balanced the equation correctly, http://www.webqc.org/balance.php feel free to plug in the reaction equation to check for yourself

OpenStudy (australopithecus):

oh I see where I went wrong

OpenStudy (australopithecus):

The conversion makes sense because the units cancel out treat them like numbers :)

OpenStudy (australopithecus):

oh I see I made a mistake balancing

OpenStudy (australopithecus):

I made a mistake balancing the equation should be 4Na + O2 = 2 Na2O I guess I misread the formula, the conversion will be different then :) I will go over soon

OpenStudy (australopithecus):

Ok I'm back, I'm actually kind of glad I made an error. This way you can see what I'm doing better :)

OpenStudy (australopithecus):

Redoing this but you might want to look at both questions to see the changes I make, that way you can understand what I'm doing better. These problems are very systematic and easy once you learn them. Na+ O2 -> Na2O To balance we need 2 more Oxygens in the product but that means we will need double the amount of product Na + O2 ->2Na2O Now we are missing 3 Na on the reactant side so we can just multiply Na by 4 4Na + O2 ->2Na2O so we use the same conversion method, notice the units cancel out (if they were numbers they would be equal to 1). Take note I included the long solution in hope you will learn something about multiplication of fractions. Note that any number can be turned into a fraction by using 1 as the denominator (which I did in the question). \[0.400mol_{Na} *\frac{2_{Na_2O}}{4_{Na}} = \frac{0.400mol_{Na}}{1} *\frac{2_{Na_2O}}{4_{Na}}= \frac{(2_{Na_2O})0.400mol_{Na}}{(1)4_{Na}} = 0.200mol_{Na_2O}\] so now we know we have 0.200mol Na2O now we just use the general formula \[moles = \frac{Grams}{Molecular Mass}\] \[0.200mol = \frac{x_{grams}}{{61.9789g/mol}}\] Isolate the x = \[\frac{0.200mol}{1} = \frac{x_{grams}}{{61.9789g/mol}}\] = \[\frac{0.200mol}{1}*\frac{61.9789g/mol}{1} = \frac{x_{grams}}{{61.9789g/mol}}*\frac{61.9789g/mol}{1}\] = \[\frac{0.200mol(61.9789g/mol)}{1(1)} = \frac{x_{grams}(61.9789g/mol)}{{61.9789g/mol}(1)}\] = \[\frac{0.200mol(61.9789g/mol)}{(1)} = \frac{x_{grams}(1)}{(1)}\] = \[0.200mol(61.9789g/mol) = x_{grams}(1)\] = \[0.200mol(61.9789g/mol) = x_{grams}\] = \[0.200*61.9789mol(\frac{g}{mol}) = x_{grams}\] = \[0.200*61.9789\frac{mol}{1}(\frac{g}{mol}) = x_{grams}\] = \[0.200*61.9789\frac{mol(g)}{(1)mol} = x_{grams}\] = \[0.200*61.9789\frac{mol(g)}{mol} = x_{grams}\] = \[0.200*61.9789g = x_{grams}\] = 0.200*61.9789g = x x = 12.4 grams Please try to follow along with my math understanding basic algebra is important and understanding how units cancel is also pretty important. If you have any further questions please ask

OpenStudy (anonymous):

okay i understand much better now! thank you! the step by step really helped!